A cone and a cylinder both have height 7 and radius 3. Find the ratio of their volumes without actually calculating them.
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)


V cone = (1/3) *pi*r^2*h
V cyl = pi*r^2*h
V cone / V cyl = 1/3

Answer by ikleyn(52779)

.

If a cone and a cylinder have equal radii and equal heights, then the ratio of their volumes
is always 1/3, independently of concrete values of the radii and the heights.